# Electrocardiogram Instruction Tuning for Report Generation

### v2

@@ 4 Benchmarking ECG Report Generation
In this section, we create a benchmark to evaluate the ECG report generation approach. Our benchmark includes two datasets (PTB-XL and MIMIC-IV-ECG) along with three ECG-related tasks, including (1) the quality of ECG report generation; (2) the zero-shot ability under domain transfer; (3) the robustness of model against perturbed signals.
4.1 Datasets
PTB-XL. The PTB-XL dataset (Wagner et al., 2020) comprises 21,837 clinical 12-lead ECGs, each lasting 10 seconds, from 18,885 patients. Each ECG is accompanied by a corresponding report. The data includes raw ECG signals and their reports, with all recordings sampled at 500Hz over 10 seconds. We divided this dataset into training, validation, and testing subsets in a 70%:10%:20% ratio, respectively. The human experts double-check all samples in the test set to ensure data quality. We reformulate the dataset into instruction data format as mentioned in Sec 2.
MIMIC-IV-ECG. The MIMIC-IV-ECG dataset (Gow et al., ) is currently the largest publicly accessible ECG dataset, containing 800,035 paired samples from 161,352 unique subjects. Similar to PTB-XL, each sample in this dataset includes a raw ECG signal and its corresponding report, with all recordings sampled at 500Hz for 10 seconds. The division of this dataset into training, validation, and testing subsets is 80%:10%:10% ratio. Likewise, we reconstruct this dataset into an ECG instruction data template.
4.2 Evaluation Metrics
We evaluate models using nine metrics: BLEU 1-4 (Papineni et al., 2002), METEOR (Banerjee & Lavie, 2005), ROUGE 1-2 and L (Lin, 2004), CIDEr-D (Vedantam et al., 2015), and BERTScore (Zhang et al., 2019). BLEU and METEOR assess the quality of machine translation, focusing on accuracy and fluency. ROUGE-L measures sentence fluency and structure, while ROUGE-1 and ROUGE-2 examine uni-gram and bi-gram overlaps. CIDEr-D evaluates the relevance and uniqueness of generated ECG reports against a candidate set, and BERTScore assesses semantic similarity to the ground truth, ensuring content accuracy.
4.3 Tasks
Quality Report Generation. The goal of this task is to evaluate the quality of report generation following ECG instruction tuning. We utilize 10% of the MIMIC-IV-ECG and PTB-XL datasets, designated for instruction tuning, as the test set. This evaluation focuses on the syntactic and semantic similarities between the generated ECG report and the ground truth, based on diverse instructions and input ECG signals. For task assessment, we employ BLEU-1 to 4, METEOR, ROUGE 1, 2, and L, CIDEr-D, and BERTScore.
Zero-shot Generalizability. To explore the generalizability of LLMs in domain transfer scenarios following ECG instruction tuning, we trained the models on 70% of the instruction data from MIMIC-IV-ECG. We then assessed the zero-shot capabilities of these models on the PTB-XL test set, which contains high-quality ground truth evaluated by clinical experts. For model evaluation, we used the metrics BLEU-3, BLEU-4, METEOR, and ROUGE-L and calculated their average.
5 Experiments and Evaluation
5.1 Experimental Settings
Models. We use nine LLMs based on the peft [^2] https://github.com/huggingface/peft library, which directly supports the LoRA (Hu et al., 2021) module, to construct the multimodal ECG report generation model described in Section 3.2. These models include GPT-Neo (Black et al., 2021), GPT-NeoX (Black et al., 2022), GPT-J (Wang & Komatsuzaki, 2021), BLOOM (Workshop et al., 2022), OPT (Zhang et al., 2022), LLaMA-1 (Touvron et al., 2023a), LLaMA-2 (Touvron et al., 2023b), Mistral (Jiang et al., 2023), and Mistral-Instruct [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1, along with two relatively small pre-trained language models (GPT2-Medium and GPT-Large (Radford et al., 2019)) as fundamental baselines.
Implementation Details. Our experiment is implemented using PyTorch 2.1, the latest transformers (Wolf et al., 2020), and the accelerate (Gugger et al., 2022) library on 4 $\times$ A100 GPUs. We utilize the official weights for all LLMs from Hugging Face (Wolf et al., 2019) for evaluation, with sizes ranging from 2.7 billion to 70 billion parameters. For models exceeding 10 billion parameters, we use DeepSpeed [^4] https://github.com/microsoft/DeepSpeed for training. All models are trained on both MIMIC-IV-ECG and PTB-XL instruction datasets for 5 epochs with a learning rate of 2e-5 and a batch size of 64. We adopt a linear optimizer scheduler with a warm-up ratio of 0.03. The best performance of each model is determined by training on the training set, selecting from the validation set, and evaluating on the test set. More details about instruction tuning can be found in the appendix.
@@ 5.2.1 Quality Evaluation
Performance on MIMIC-IV-ECG. Table 1 and 3 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG. The results show that all LLMs exhibit superior performance compared to smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics. Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics. For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore.
The observed performance underscores the adeptness of LLMs in generalizing from signal data, showcasing enhanced proficiency in aligning ECG signal representations with corresponding textual information. This highlights the significant potential of LLMs in the domain of medical signal-to-text generation. Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in the majority of evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
@@
Performance on PTB-XL. As shown in Table 2, the models exhibit reduced performance on PTB-XL compared to MIMIC-IV-ECG, which is attributable to the smaller scale of the instruction data in PTB-XL. This underscores the importance of data scale in enhancing instruction-based ECG report generation. Moreover, similar to the MIMIC-IV-ECG results, all LLM-based models show significant improvement over SLMs. Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score. The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics, because of their strong general instruction-following capabilities. Consequently, we select LLaMA-2 and Mistral-Instruct as the default architectures for our framework.
5.2.2 Zero-shot Evaluation in Domain Transfer.
In Figure 3, we present the evaluation of the zero-shot learning capabilities of various LLMs, which is trained on the MIMIC-IV-ECG dataset and then tested on PTB-XL (unseen dataset). The assessed models include BLOOM, OPT, LLaMA-1, and Mistral. Firstly, all selected LLMs undergo instruction tuning on the MIMIC-IV-ECG train set, followed by zero-shot testing on the PTB-XL test set verified by human experts, denoted as Zero-shot IT. We also measure the performance of each model in report generation without prior ECG-specific instruction tuning, denoted as Zero-shot W/O IT. PTB-XL IT represents training on the PTB-XL train set and then evaluated on the PTB-XL test set. Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the models’ ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT. The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
5.2.3 Robust Analysis with Perturbed ECG Signal.
In a noise stress evaluation (Wang et al., 2019), we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness. Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG training set and tested on corresponding noise-added signals from its test set. The results, illustrated in Figure 4, show a performance decline in all LLMs as SNR increased, highlighting the significant interference of ECG noise. Furthermore, as shown in Table 1, Mistral also excelled in tests on noise-free datasets, suggesting a synergistic effect between clean and noisy test sets. The results demonstrate Mistral’s strong resistance to perturbations. Even with more severe noise, it maintained robustness in terms of ROUGE-L and METEOR metrics. Developing an even more robust framework is a goal for future research.
@@ 5.3 Analysis
5.3.1 Instruction Tuning Visualization.
Figure 6 compares the convergence curves of the instruction tuning loss and the METEOR score between GPT-Neo (2.7B), BLOOM (7B), OPT (6.7B), and LLaMA-2 (7B) on the MIMIC-IV-ECG train and validation datasets. We observe that larger models with more parameters can converge to a smaller loss and achieve higher performance on the METEOR score. Notably, an increase in model size correlates with higher performance and lower loss, suggesting that larger models have the potential for better performance.
@@
5.3.2 Scalability Analysis.
To investigate whether ECG instruction tuning on larger-scale models yields better results, we validated LLaMA-2 models of 7B and 13B parameter sizes on both MIMIC-IV-ECG and PTB-XL datasets. As depicted in 7, an upward trend in all evaluation metrics is observed with a gradual increase in model size. However, it is noteworthy that the gains in performance associated with increasing model size are not particularly significant. For example, the F-1 score for the 70B model on the PTB-XL dataset exhibits a marginal increase of 0.02 over the 13B model. Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model. Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance (Wei et al., 2022).
5.3.3 Ablation Study on ECG Instruction Tuning.
To validate the effectiveness of adopting an instruction tuning paradigm for aligning ECG signals and report representations, we conducted the following ablation experiment. We selected BLOOM, OPT, LLaMA-1, and Mistral as LLM backbones and removed the instruction tuning strategy, allowing the language models to fine-tune directly in the setting of generating text from ECG signals only. As shown in Figure 5, aligning ECG and report representations using direct fine-tuning, as opposed to instruction tuning, results in a significant decline across all metrics. Particularly in the Mistral model, the instruction tuning paradigm notably outperforms direct fine-tuning. These results demonstrate that, compared to fine-tuning, instruction tuning more effectively enhances the generalization ability of LLMs on unseen tasks/data (Ouyang et al., 2022).
@@
5.3.4 Qualitative Results.
In Figure 8, we randomly select two samples generated by MEIT using LLaMA-2 and Mistral-Instruct as the LLM backbones. The consistent key information, highlighted in blue, indicates that both models have successfully learned important patterns from the ECG signals. Overall, the models’ results align with the ground truth, accurately identifying cardiac abnormalities from the ECG signals. Furthermore, both models provide detailed explanations of abnormal ECG signal details, such as ‘ischemia’ from sample 1 and ‘right bundle branch block’ from sample 2. These generated reports demonstrate the efficacy of our method.
6 Related Work
Instruction Tuning. Instruction tuning (Zhang et al., 2023a; Wang et al., 2023) is designed to enhance the zero-shot capabilities of LLMs on new tasks using either machine-generated or human-annotated instructions. Notable examples such as InstructGPT (Ouyang et al., 2022), FLAN-PaLM (Chung et al., 2022), and Alpaca (Taori et al., 2023) utilize various methods, including manual collection and human feedback, to generate instruction template data for fine-tuning. In the realm of multimodal learning, models like LLaVA (Liu et al., 2023b), MiniGPT-4 (Zhu et al., 2023), and AnyMAL (Moon et al., 2023) demonstrate the utility of multimodal instruction data to enhance learning across different modalities. However, medical signals such as ECG and their corresponding text remain underexplored. Our work bridges this gap by developing a specialized ECG-related instruction tuning framework, thereby improving the generation of ECG reports.
Medical Report Generation. Different from medical vision-language alignment (Wan et al., 2023a), medical Report Generation can be categorized into three main approaches: (1) Template Selection and Generation, where methods like HRGR (Li et al., 2018) focus on sentence creation or template retrieval, and CMAS (Jing et al., 2017) uses dual writers for differentiating regions; (2) Data Integration and Coherence, illustrated by PPKED’s (Liu et al., 2021b) knowledge merging for consistency, and CA’s contrasting of abnormal features (Ma et al., 2021); (3) Cross-Modal Alignment, with (Chen et al., 2022b; Qin & Song, 2022) using attention mechanisms for text-visual feature correlation. Recent advancements (Gu et al., 2024; Tanida et al., 2023; Li et al., 2023c) focus on local visual feature learning and multi-level cross-modal alignment, showcasing the evolution of semantic matching in radiology report generation. However, these approaches mainly focus on radiology report generation, directly applying these approaches to ECG data is challenging because the patterns between medical images and ECG biomedical signals are significantly different.
Signal-to-Text Generation. Signal-to-Text (S2T) generation, predominantly speech-to-text, converts speech signals into text. Extensive research (Baevski et al., 2020; Chen et al., 2020; Liu et al., 2021a; Zheng et al., 2021; Hsu et al., 2021; Bapna et al., 2021) has extensively studied patterns in speech signals, analyzing both temporal and spectral aspects. Although the field of ECG report generation is largely unexplored, the use of speech signal research techniques in the creation of ECG reports presents significant challenges due to fundamental differences: (1) Speech signals for S2T tasks typically have a sampling frequency of up to 16,000 - 60,000 Hz (Lavry, 2012; Painter & Spanias, 2000; Byrne et al., 1994) and are monochannel, while ECG signals usually have a lower sampling frequency of around 500 Hz and encompass 12 channels. (2) Speech signals can be discretized into acoustic units corresponding to specific words (Défossez et al., 2022), but ECG signals lack such discrete units. Certain cardiac conditions, like bundle branch block, can obscure or distort the heartbeats patterns in ECG readings. (3) Speech signals and their transcriptions maintain strong temporal alignment, as each word correlates with a specific acoustic pattern, and the transcription follows the temporal sequence of speech (Zhang et al., 2023b). In contrast, ECG reports evaluate the cardiac patterns on a global level, assessing the entire ECG period rather than focusing on individual acoustic-like units in speech. These fundamental differences between speech-text pairs and ECG records, in both signal and text terms, highlight the need for S2T methods specifically designed for ECG data. Furthermore, the scarcity of research in the generation of ECG reports highlights an urgent need for comprehensive benchmarks using large-scale, publicly accessible data sets.
@@ 7 Conclusion and Future Work
Broader Impact
Our MEIT framework first attempt to addresses automatic ECG report generation through ECG instruction tuning, establishing the first comprehensive benchmark for this process with mainstream LLMs. In this work, we mainly focus on generating reports from multimodal ECG instructions using LLMs. However, the generated results are not fully explainable or controllable, even though the generation procedure is transparent and trackable. This is because the underlying theory of LLMs remains largely unexplored, necessitating further investigation to ensure the quality and safety of the generated content. In the future, we aim to enable LLMs to utilize external, expert-verified knowledge databases, such as clinical protocols and medical textbooks, to enhance the explainability of the generated ECG reports.
Appendix A Hyper-parameters of ECG Instruction Tuning

### v3

@@ 5.1 Experimental Settings
two relatively small pre-trained language models (GPT2-Medium and GPT-Large [29]) as fundamental baselines.
Implementation Details. Our study utilized PyTorch 2.1, transformers [30], and accelerated on A100
@@
GPUs with LLMs from Hugging Face [31] ranging from 2.7 to 70 billion parameters. For larger models, we used DeepSpeed [^3] https://github.com/microsoft/DeepSpeed. The training covered 5 epochs on MIMIC-IV-ECG and PTB-XL with a 2e-5 learning rate and 64 batch size, employing a linear optimizer with a 0.03 warm-up ratio. For text preprocessing, we initially remove all instances of the ‘nan’ string and sentences that consist solely of numerical values. Subsequently, we discard any samples whose reports contain fewer than 5 tokens. For, ECG encoder, we adopt random initialization. Additionally, the default number of generated prompts from GPT-4 is 256, more training, visualization details about ECG instruction tuning are illustrated in Appendix A.2, Section  5.3, and Appendix A.4.
@@ 5.2.1 Quality Evaluation
Performance on MIMIC-I V-ECG. Table 3 and 1 present the results of various types of language encoders $\mathcal{F}_{l}(\cdot)$ on MIMIC-IV-ECG. The results show that all LLMs perform better than smaller language models (SLMs), such as GPT2-Medium and GPT2-Large, across all evaluation metrics. Notably, from GPT-Neo to Mistral-Instruct, LLM-based backbones achieve a significant margin over SLMs in all metrics. For instance, compared to GPT2-Large, the METEOR score increases in the range of 0.132 to 0.18 from GPT-Neo to LLaMA-2, and Mistral-Instruct outperforms GPT2-Large with an improvement of 0.18 in the ROUGE-L score and 0.134 in the F-1 of BERTScore.
@@
The observed performance underscores the adeptness of LLMs in generalizing from signal data, showcasing enhanced proficiency in aligning ECG signal representations with corresponding textual information. This highlights the significant potential of LLMs in medical signal-to-text generation. Particularly, LLaMA-2 and Mistral-Instruct surpass their counterparts in most evaluative metrics, suggesting that models pre-tuned with general instructions are more adept at learning ECG-text alignment.
Performance on PTB-XL. As shown in Table 2, the models exhibit reduced performance on PTB-XL compared to MIMIC-IV-ECG, which is attributable to the smaller scale of the instruction data in PTB-XL. This underscores the importance of data scale in enhancing instruction-based ECG report generation. Moreover, similar to the MIMIC-IV-ECG results, all LLM-based models show significant improvement over SLMs. Specifically, LLaMA-2 surpasses GPT2-Large by 0.134 in the BLEU-3 metric, while LLaMA-1 achieves a 0.103 improvement in the METEOR score. The overall experimental results reveal that Mistral-Instruct and LLaMA-2 are consistently the top two performers across most metrics because of their strong general instruction-following capabilities.
5.2.2 Zero-shot Evaluation in Domain Transfer.
In Figure 3, we present the evaluation of the zero-shot learning capabilities of various LLMs, which is trained on the MIMIC-IV-ECG dataset and then tested on PTB-XL (unseen dataset). The assessed models include BLOOM, OPT, LLaMA-1, and Mistral. Firstly, all selected LLMs undergo instruction tuning on the MIMIC-IV-ECG train set, followed by zero-shot testing on the PTB-XL test set verified by human experts, denoted as Zero-shot IT. We also measure the performance of each model in report generation without prior ECG-specific instruction tuning, denoted as Zero-shot W/O IT. PTB-XL IT represents training on the PTB-XL train set and then evaluated on the PTB-XL test set. Notably, although Zero-shot IT shows a slight degradation compared to PTB-XL IT, the results still indicate a variance in the model’s ability to generalize to an unseen dataset with instruction tuning (IT), compared to Zero-shot W/O IT. The involvement of ECG instruction tuning on MIMIC-IV-ECG enables the models to achieve superior zero-shot performance on the unseen PTB-XL dataset, indicating the necessity of instruction tuning in enhancing the models’ zero-shot ability on unseen datasets in ECG report generation.
5.2.3 Robust Analysis with Perturbed ECG Signal.
In a noise stress evaluation [32], we added Gaussian noise to ECG signals at signal-to-noise ratios (SNRs) of 0.05, 0.1, 0.15, and 0.2 during testing to assess model robustness. Our experiments utilized four LLM architectures: BLOOM, OPT, LLaMA-1, and Mistral, each trained on clean ECG signals from the MIMIC-IV-ECG
@@
training set and tested on corresponding noise-added signals from its test set. The results, illustrated in Figure 4, show a performance decline in all LLMs as SNR decreased, highlighting the significant interference of ECG noise. Furthermore, as shown in Table 1, Mistral also excelled in tests on noise-free datasets, suggesting a synergistic effect between clean and noisy test sets. The results demonstrate Mistral’s strong resistance to perturbations. Even with more severe noise, it maintained robustness regarding ROUGE-L and METEOR metrics. Developing an even more robust framework is a goal for future research.
Table 4: Performance comparison of the proposed concatenated-fusion method and other mainstream fusion variants. We evaluate these methods on the MIMIC-IV-ECG dataset, using BLEU-4, METEOR, ROUGE-L, and CIDEr-D metrics. We take LLaMA-1 7B as the LLM backbone here. heavy teal color indicates the highest results.
Framework Method BLEU-4 METEOR ROUGE-L CIDEr-D
LLaVA Straightforward input 0.529 0.737 0.712 4.99
Flamingo Trainable cross-attention 0.527 0.768 0.715 5.11
MEIT Concatenated-fusion 0.543 0.761 0.724 5.26
@@ 5.3 Analysis
Instruction Tuning Visualization. Figure 5 compares the convergence curves of the instruction tuning loss and the METEOR score between GPT-Neo (2.7B), BLOOM (7B), OPT (6.7B), and LLaMA-2 (7B) on the MIMIC-IV-ECG train and validation datasets. We observe that larger models with more parameters can converge to a more minor loss and achieve higher performance on the METEOR score. Notably, an increase in model size correlates with higher performance and lower loss, suggesting that larger models have the potential for better performance.
Analysis of ECG Modality Alignment. To study the effectiveness of our proposed concatenated-fusion method for ECG modality alignment, we compare it with other fusion approaches such as direct input in LLaVA [33] and additional trainable cross-attention layer in Flamingo [12]. For straightforward input, we follow the design of LLaVA by directly concatenating the ECG encoder’s output embeddings with the sentence’s embeddings before inputting them into the LLM backbones. For the second comparison method, we follow Flamingo by adding a trainable cross-attention layer within the attention block. From Table 4, we observe that the Concatenated-fusion method outperforms the trainable cross-attention method of Flamingo in most metrics and is consistently superior to the Straightforward input method of LLaVA. Consequently, the concatenated fusion is more effective for the LLM backbone’s alignment with fine-grained ECG patterns without necessitating additional trainable parameters.
@@
Scalability Analysis. To investigate whether ECG instruction tuning on larger-scale models yields better results, we validated LLaMA-2 models of 7B, 13B, and 70B parameter sizes on both MIMIC-IV-ECG and PTB-XL datasets. As depicted in Figure 6, an upward trend in all evaluation metrics is observed with a gradual increase in model size.
However, it is noteworthy that the gains in performance associated with increasing model size are not particularly significant. For example, the F-1 score for the 70B model on the PTB-XL dataset exhibits a marginal increase of 0.02 over the 13B model. Similarly, on the MIMIC-IV-ECG dataset, the 70B model’s F-1 score is only 0.01 higher than that of the 13B model. Therefore, we conjecture that enhancing both data scale and model size concurrently is necessary to achieve superior performance [34].
Ablation Study on ECG Instruction Tuning. We conducted an ablation study to evaluate instruction tuning’s impact on aligning ECG signals with report representations. Utilizing LLMs such as BLOOM, OPT, LLaMA-1, and Mistral without instruction tuning, we allowed direct learning from ECG signals. The findings, illustrated in Figure 7, indicate a significant performance drop across all metrics without instruction tuning, particularly in Mistral. This underscores instruction tuning’s superiority in enhancing LLMs’ generalization to new tasks/data over direct fine-tuning [35].
Qualitative Results. In Figure 8, we randomly select two samples generated by MEIT using LLaMA-2 and Mistral-Instruct as the LLM backbones. The consistent key information, highlighted in blue, indicates that both models have successfully learned important patterns from the ECG signals. Overall, the models’ results align with the ground truth, accurately identifying cardiac abnormalities from the ECG signals. Furthermore, both models provide detailed explanations of abnormal ECG signal details, such as ‘ischemia’ from sample 1 and ‘right bundle branch block’ from sample 2. These generated reports demonstrate the efficacy of our method.
6 Related Works
Instruction Tuning. Instruction tuning [36, 37] boosts zero-shot learning in LLMs for new tasks using instructions. Notable models like InstructGPT [35], FLAN-PaLM [38],
@@ 6 Related Works
and Alpaca [39] fine-tune with instruction data through various methods, including human feedback. Similarly, multimodal models such as LLaVA [33], MiniGPT-4 [40], and AnyMAL [41] benefit from multimodal instructions for enhanced learning. However, instruction tuning for medical signals like ECG remains underexplored. Our research develops a targeted instruction-tuning framework for ECG report generation, addressing this gap.
Medical Report Generation. Medical report generation [42]involves three strategies: (1) Template Selection and Generation, highlighted by HRGR [43] and CMAS [44]; (2) Data Integration and Coherence, as seen in PPKED [45] and CA [46]; (3) Cross-Modal Alignment, with efforts like [47, 48]. While recent research has advanced these areas, applying them to ECG data is challenging due to the distinct nature of ECG signals compared to medical images.
LLM for ECG Understanding. Only a few research efforts have focused on utilizing LLMs for ECG-related tasks [49, 50, 51]. In the studies by [49, 51], the authors manually convert ECG signals into text features before feeding them into LLMs, rather than using the original ECG data. This approach overlooks the modality patterns inherent in the signals. Moreover, these studies only employ LLMs for classifying diseases from ECG data and do not explore the generation of medical reports. In contrast, [50] attempts ECG report generation by using handcrafted ECG features as inputs instead of the original signals. However, they have not publicly made any code or models available, making direct comparisons challenging.
@@ 7 Conclusion and Future Work
Appendix A Appendix.
A.1 Limitations and Future Work
In this work, we propose an effective and efficient method for generating ECG reports. Since this is a generative task, the content produced may occasionally lead to hallucinations. This is particularly pertinent in medical contexts such as ECG report generation, where accuracy is of utmost importance. To mitigate this issue, future iterations of our work could incorporate retrieval-augmented methods [14], information optimized strategies [52, 53, 54]. For instance, by integrating a retrieval mechanism that accesses a database of verified medical or biological knowledge [55] or previous ECG reports, the model could be guided towards generating more accurate and reliable outputs.
A.2 Hyper-parameters of ECG Instruction Tuning